# Capacity Approximations for Insertion Channels with Small Insertion Probabilities

### v3

@@ IV-A Calculation of $H(A^{n},B^{n})$
Note that in (74), the optimal distribution of the input run lengths $p_{L(2)}$ is used. To establish a bound for this term, as in [23, Lemma V.5], let us assume $p^{*}_{L(k)}(l_{1},\dots,l_{k})=2^{-\sum_{i=1}^{k}l_{i}}$. Consider any positive integer $k$, $\kappa<\infty$, and $H(\mathbb{X})>1-\alpha^{\gamma}$ with $\gamma>1/2$. Then, [23, Lemma V.5] implies
$\mid p_{L(k)}(l_{1},\dots,l_{k})-p^{*}_{L(k)}(l_{1},\dots,l_{k})\mid\leq\kappa\sqrt{k}\alpha^{\gamma/2}.$ (76)
Inserting (76) to the first term of the (74), one can write
$H(\check{K}|X^{n},Y)=A_{1}\alpha-\epsilon_{2}+\epsilon_{3},$ (77)
@@
$A_{1}=\frac{1}{2}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}$ $(r_{j+1}+1)2^{-r_{j}}2^{-r_{j+1}}h\left(\frac{1}{r_{j+1}+1}\right),$ (78)
$-\alpha^{1+\gamma/2-\epsilon}\leq\epsilon_{3}\leq\alpha^{1+\gamma/2-\epsilon},$ (79)
@@
For a simple insertion channel model with a stationary and ergodic input process $\mathbb{X}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$; the difference between $H(K|X^{n},Y)$ and $H(\check{K}|X^{n},\check{Y})$, corresponding to the original insertion and perturbed insertion processes, respectively, satisfies:
$|H(K|X^{n},Y)-H(\check{K}|X^{n},\check{Y})|\leq 4\alpha^{1+\gamma-\epsilon/2},$ (80)
@@
The proof strategy mirrors that of [23, Lemma V.18]. We utilize the perturbed insertion process $\check{Z}$ and define $U(X^{n},A^{n},Z^{n})\in\{t,0,1\}^{|{Y}|}$ for each bit in the output of size $|Y|$. Sequentially, for each bit in $Y$, we define the elements of $U$ based on the corresponding bit in $\check{Y}$ and the perturbed insertion process as follows:
@@
$(X^{n},Y,K)$–→$\mathrel{\mathop{\leftarrow\joinrel\relbar\joinrel\relbar\joinrel\rightarrow}\limits^{(U,Z)}}(X^{n},\check{Y},\check{K}).$ (81)
@@
With the perturbed process, we further define $\check{z}=P(\check{Z}_{j}=1)$ for each $j$. The number of insertions reversed in a random run is at most
$\alpha^{2}\sum_{l_{0},l_{1}}p_{{L}}(l_{0},l_{1})(l_{0}+l_{1})^{2},$ (83)
in expectation, which can be bounded above by $4\alpha^{2}\mathbb{E}[{L}^{2}]$.
As in [23, Lemma V.3], there exists $\alpha_{0}$ such that the following occurs: Consider any $\gamma>1/2$, and define $\ell\equiv\lfloor 2\gamma\log(1/\alpha)\rfloor$. For all $\alpha<\alpha_{0}$, if $\mathbb{X}\in\mathcal{S}$ is such that $H(\mathbb{X})>1-\alpha^{\gamma}$, we have
@@
Using [23, Lemma V.3] and ${L}\leq 1/\alpha$ w.p. 1, we have
@@
which implies
@@
$\qquad\cdot 2^{-l_{j}}2^{-l_{j+1}}h\big{(}\frac{1}{l_{j+1}+1}\big{)}\bigg{)}+\zeta,$ (90)
@@
$-\alpha^{1+\gamma/2-\epsilon}-4\alpha^{1+\gamma-\epsilon/2}-\frac{\alpha^{2}}{2}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1})(r_{j+1}+1)2^{-r_{j}}2^{-r_{j+1}}h\left(\frac{1}{r_{j+1}+1}\right)\leq\epsilon_{1}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\leq 4\alpha^{1+\gamma-\epsilon/2}+\alpha^{1+\gamma/2-\epsilon},$
$-2h(z,v)\leq\eta\leq\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+2h(z,v),$
@@ IV-E Achiveability
$C(\alpha)=1+\alpha\log(\alpha)+G_{1}\alpha+\mathcal{O}(\alpha^{3/2-\epsilon}).$ (93)
@@
Since $\mathbb{X}$ is i.i.d. Bernoulli(1/2) with size $n$, it has a run-length distribution $p_{L}(l)=2^{-l}$. Additionally, the corresponding output $Y(X^{*,n})$ is also i.i.d. Bernoulli with length $L=n+\text{Bin}(n,\alpha)$. Hence, we have
$H(Y,L)$ $=H(L)+H(Y|L)$
$=H(L)+\sum_{l=n}^{2n}H(Y|L=l)p_{L}(l)$
$=H(L)+\sum_{l=n}^{2n}l2^{-l}$
$=\mathcal{O}(\log(n))+n(1+\alpha).$ (94)
As a result,
$H(Y(X^{*,n}))$ $=\mathcal{O}(\log(n))+n(1+\alpha).$ (95)
@@ IV-F Converse
We further define $F=F(\mathbb{X},\mathbb{D})$ as a binary vector with the same length as the channel output. The elements of $F$ are 1 wherever the corresponding bit in $Y_{L^{*}}$ is flipped relative to $Y$, and 0 otherwise. The entropy $H(F)$ can be calculated by:
@@
where $L_{F}$ is the length of $F$. The rest of the proof follows the same steps as [22, Lemma III.2]. ∎
@@
For any $\epsilon>0$ there exists there exists $\alpha_{0}=\alpha_{0}(\epsilon)>0$ such that the following happens. For any $L^{*}\in\mathbb{N}$ and any $\mathbb{X}\in\mathcal{S}_{L^{*}}$ if $\alpha<\alpha_{0}(\epsilon)$, then
$I(\mathbb{X})\leq 1+\alpha\log\alpha+G_{1}\alpha+\alpha^{2-\epsilon}(1+\alpha^{1/2}L^{*}).$ (102)
@@
We can write
@@
since $Y(X^{n})$ contains $n+\text{Binomial}(n,\alpha)$ bits. The proof involves using the lower bound in Corollary 1 along with $h(z,v)$. Note that for $H(\mathbf{X})>1+2\alpha\log\alpha$, it is also true that $H(\mathbf{X})>1-\alpha^{1-\epsilon}$. Hence Corollary 1 with $\alpha=1-\epsilon$ applies. Furthermore, recall that the joint PMF of $Z$ and $V$ was given in (36). One can upper bound the probabilities as
$P(z=0,v=0)$ $\leq 1-2\alpha^{2}\mathbb{E}[L_{0}],$ (104)
$P(z=1,v=0)$ $\leq\alpha^{2}\mathbb{E}[L_{0}]\leq 2\alpha^{2}\mathbb{E}[L_{0}],$
$P(z=1,v=1)$ $\leq\alpha^{2}\mathbb{E}[L_{0}]\leq 2\alpha^{2}\mathbb{E}[L_{0}].$
We employ [22, Lemma IV.3], which gives $\mathbb{E}[L_{0}]\leq K_{1}(1+\sqrt{\alpha(\log(1/\alpha))^{3}L^{*}})$; hence, we obtain
$h(z,v)\leq 0.5\alpha^{2-\epsilon}(2+0.5\alpha^{1/2}L^{*}),$ (105)
for all $\alpha\leq\alpha_{0}$ with $\alpha_{0}=\alpha_{0}(\epsilon)>0$. Also noting
@@ V Capacity Approximation for the Gallager Insertion Channel
Let $C(\alpha)$ be the capacity of the insertion channel with insertion probability $\alpha$. Then, for small $\alpha$ and any $\epsilon>0$, we have
$C(\alpha)=1+\alpha\log(\alpha)+G_{2}\alpha+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{O}(\alpha^{3/2-\epsilon})}},$ (107)
@@
$+{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\frac{1}{4}}}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}(a+b+2)2^{-a}2^{-b}h\left(\frac{a+1}{a+b+2}\right),$ (108)
@@ V-A Calculation of $H(A^{n},B^{n})$
$P(a,b^{1},b^{2})=\begin{cases}\frac{\alpha}{4},&\text{if }(a,b^{1},b^{2})\in\{(1,1,0),(1,0,1),\\ &\qquad\qquad(1,1,1),(1,0,0)\}\\ 1-\alpha,&\text{if }(a,b^{1},b^{2})=(0,0,0)\\ 0,&\text{otherwise}\end{cases}$
@@
As in Section IV-B, we introduce a modified insertion process for the Gallager insertion channel by defining $\hat{A}^{n}$ and $\hat{B}^{n}=(\hat{B}^{1,n},\hat{B}^{2,n})$. For the extended runs, the modified process keeps the runs with at most one insertion as they are while reversing more insertions.
@@
$P$ $(z=a,v^{1}=b,v^{2}=c)=\begin{cases}&1-\alpha+\mathbb{E}[\alpha(1-\alpha)^{L_{0}+1}],\\ &\qquad\text{if }(a,b,c)=(0,0,0)\\ &\frac{1}{4}\left(\alpha-\mathbb{E}[\alpha(1-\alpha)^{L_{0}+1}]\right),\\ &\qquad\text{if }(a,b,c)\in\mathcal{A}\\ &0,\qquad\text{otherwise},\end{cases}$ (118)
@@
As in Sec. IV-B, one can use (118) to calculate $H(Z^{n},{V}^{1,n},{V}^{2,n})\leq nh(z,v^{1},v^{2})$ and show that it does not affect lower order terms. Comparing (118) with (36) (from the previous channel model), one can easily verify that $h(z,v^{1},v^{2})\leq h(z,v)$.
@@
Consider the Gallager insertion channel model with any $\mathbb{X}$ such that $\mathbb{E}[L_{0}\log L_{0}<\infty]$. Then, we have
@@
where $-2h(z,v^{1},v^{2})\leq\eta\leq\frac{1}{2}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+\frac{1}{2}\alpha^{2}(\mathbb{E}[L_{0}]-1)+2h(z,v^{1},v^{2})$.
@@
where $0\leq\delta\leq\frac{1}{2}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+\frac{1}{2}\alpha^{2}(\mathbb{E}[L_{0}]-1)$.
@@
    No insertion: In this case, there is no ambiguity and no contribution to $H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})$ as discussed in Section IV-B.
@@
    Replace a bit with ‘11’: Inserting ‘11’ into the run of ‘0’s reveals all information regarding $\hat{D}$, $\hat{B}^{1,n}$, and $\hat{B}^{2,n}$ given the marker vector $\hat{K}$. Hence, there’is no contribution to $H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})$.
@@
Hence, considering the insertions with ambiguity (and the intersecting events of insertions with ‘01’ and ‘10’), the overall contribution to $H(\hat{A}^{n},\hat{B}^{n}|x^{n},\hat{y},\hat{k})$ can be written as:
@@
$P(\hat{A}_{i}=1)\in(\alpha-2l_{(i)}\alpha^{2},\alpha).$ (131)
@@
where $0\leq\delta\leq\frac{1}{2}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+\frac{1}{2}\alpha^{2}(\mathbb{E}[L_{0}]-1)$, which completes the proof of Lemma 16. ∎
@@
For the Gallager insertion channel with a stationary and ergodic input process $\mathbb{X}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$, we have:
$H(K|X^{n},Y)=$ $\frac{\alpha}{4}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}(a+b+2)2^{-a}2^{-b}h\left(\frac{a+1}{a+b+2}\right)+\epsilon_{1},$ (133)
@@
$-\alpha^{1+\gamma/2-\epsilon}-4\alpha^{1+\gamma-\epsilon/2}-\frac{\alpha^{2}}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1})(r_{j}+r_{j+1}+2)h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right)\leq\epsilon_{1}\leq 4\alpha^{1+\gamma-\epsilon/2}+\alpha^{1+\gamma/2-\epsilon},$ (134)
@@
To compute $H(K|X^{n},Y)$, we define the perturbed insertion process $\check{\mathbb{Z}}$ as in Section IV-C, and follow the same steps. We use the same definitions $\check{K}$, $\check{Y}$, and $\bar{Y}$. With this process, we assume that there is at most one insertion on consecutive two runs.
@@
For the Gallager insertion channel with the perturbed insertion process and a stationary and ergodic input process $\mathbb{X}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$, we have:
$H(\check{K}|X^{n},\check{Y})=$ $\frac{\alpha}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1}+2)2^{-r_{j}}2^{-r_{j+1}}h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right)-\epsilon_{2}+\epsilon_{3}$ (135)
@@
$-\alpha^{1+\gamma/2-\epsilon}\leq\epsilon_{3}\leq\alpha^{1+\gamma/2-\epsilon}.$ (137)

### v4

@@ IV-A Calculation of $H(A^{n},B^{n})$
Note that in (72), the optimal distribution of the input run lengths $p_{L(2)}$ is used. To establish a bound for this term, as in [kanoria2013optimal, Lemma V.5], let us assume $p^{*}_{L(k)}(l_{1},\dots,l_{k})=2^{-\sum_{i=1}^{k}l_{i}}$. Consider any positive integer $k$, $\kappa<\infty$, and $H(\mathbb{X})>1-\alpha^{\gamma}$ with $\gamma>1/2$. From [kanoria2013optimal, Lemma V.5], we have
$\sum_{l_{1}=1}^{\infty}\cdots\sum_{l_{k}=1}^{\infty}\left|p_{L(k)}(l_{1},\dots,l_{k})-p^{*}_{L(k)}(l_{1},\dots,l_{k})\right|\leq\kappa\sqrt{k}\alpha^{\gamma/2},$ (74)
which implies that each term inside the summation also satisfies the following inequality, as all terms are nonnegative:
$\left|p_{L(k)}(l_{1},\dots,l_{k})-p^{*}_{L(k)}(l_{1},\dots,l_{k})\right|\leq\kappa\sqrt{k}\alpha^{\gamma/2}.$ (75)
With $\mathbb{X}\in\mathcal{S}_{\lfloor 1/\alpha\rfloor}$, for the summation term in (72), we define $l\equiv\lfloor 4\log(1/\alpha)\rfloor)$. Note that we choose $\ell=\lfloor 4\log(1/\alpha)\rfloor$ because, under the constraint $H(X)>1-\alpha^{\gamma}$ with $\gamma>1/2$, the run-length distribution has an exponential tail as stated in [kanoria2013optimal, Lemma V.3]. Consequently, terms involving runs longer than $\ell$ contribute only to higher-order terms and are therefore negligible in the overall asymptotic expansion.
We have
$\frac{\alpha}{2}\sum_{r_{j}=1}^{l}$ $\sum_{r_{j+1}=1}^{l}(r_{j+1}+1)p_{L(2)}(r_{j},r_{j+1})h\left(\frac{1}{r_{j+1}+1}\right)$
$=$ $\frac{\alpha}{2}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j+1}+1)2^{-r_{j}-r_{j+1}}h\left(\frac{1}{r_{j+1}+1}\right)-\epsilon_{3},$ (76)
resulting in
$\lim_{n\rightarrow\infty}\frac{1}{n}H(\check{K}|X^{n},Y)=A_{1}\alpha-\epsilon_{2}+\epsilon_{3},$ (77)
@@
$A_{1}$ $=\frac{1}{2}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j+1}+1)2^{-r_{j}}2^{-r_{j+1}}h\left(\frac{1}{r_{j+1}+1}\right)$
$=\frac{1}{2}\sum_{r_{j+1}=1}^{\infty}(r_{j+1}+1)2^{-r_{j+1}}h\left(\frac{1}{r_{j+1}+1}\right),$ (78)
$\alpha^{1+\gamma/2-\epsilon}\leq\epsilon_{3}\leq\alpha^{1+\gamma/2-\epsilon},$ (79)
@@
For a simple insertion channel model with a stationary and ergodic input process $\mathbb{X}\in\mathcal{S}_{\lfloor 1/\alpha\rfloor}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$, and for any $\epsilon>0$, the difference between $H(K|X^{n},Y)$ and $H(\check{K}|X^{n},\check{Y})$, corresponding to the original insertion and perturbed insertion processes, respectively, satisfies:
$\lim_{n\rightarrow\infty}\frac{1}{n}|H(K|X^{n},Y)-H(\check{K}|X^{n},\check{Y})|\leq 4\alpha^{1+\gamma-\epsilon/2}.$ (80)
@@
The proof strategy mirrors that of [kanoria2013optimal, Lemma V.18]. We utilize the perturbed insertion process $\check{Z}$ and define $U(X^{n},A^{n},Z^{n})\in\{t,0,1\}^{|{Y}|}$ for each bit in the output of size $|Y|$. Sequentially, for each bit in $Y$, we define the elements of $U$ based on the corresponding bit in $\check{Y}$ and the perturbed insertion process as follows:
@@
$(X^{n},Y,K)$ $\mathrel{\mathop{\leftarrow\joinrel\relbar\joinrel\relbar\joinrel\rightarrow}\limits^{(U,Z)}}(X^{n},\check{Y},\check{K}).$ (81)
@@
With the perturbed process, we further define $\check{z}=P(\check{Z}_{j}=1)$ for each $j$, which is the probability of having a reversed insertion for the $j$-th index. The number of insertions reversed in a random run is at most
$\alpha^{2}\sum_{l_{0},l_{1}}p_{{L(2)}}(l_{0},l_{1})(l_{0}+l_{1})^{2},$ (83)
in expectation, where $p_{(2)}(l_{0},l_{2})$ is the joint distribution of input run lengths for $2$ runs. This expression can be bounded above by $4\alpha^{2}\mathbb{E}[{L}^{2}]$.
As in [kanoria2013optimal, Lemma V.3], there exists $\alpha_{0}$ such that the following occurs: Consider any $\gamma>1/2$, and define $\ell\equiv\lfloor 2\gamma\log(1/\alpha)\rfloor$. For all $\alpha<\alpha_{0}$, if $\mathbb{X}\in\mathcal{S}$ is such that $H(\mathbb{X})>1-\alpha^{\gamma}$, we have
@@
Using [kanoria2013optimal, Lemma V.3] and ${L}\leq 1/\alpha$ with probability 1, we have
@@
which implies the number of reversed insertion in a run can be upper bounded by $80\alpha^{1+\gamma}$, i.e., we have
$\alpha^{2}\sum_{l_{0},l_{1}}p_{{L(2)}}(l_{0},l_{1})(l_{0}+l_{1})^{2}\leq 80\alpha^{1+\gamma}.$ (86)
Furthermore, since the length of a run is at least 1, normalizing (86) by 1 yields a bound for $\check{z}$ as well, which is
@@
$\hskip 18.49988pt\cdot 2^{-l_{j}}2^{-l_{j+1}}h\big(\frac{1}{l_{j+1}+1}\big)\bigg)+\zeta,$ (91)
@@
$-\alpha^{1+\gamma/2-\epsilon}-4\alpha^{1+\gamma-\epsilon/2}-\frac{\alpha^{2}}{2}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1})(r_{j+1}+1)h\left(\frac{1}{r_{j+1}+1}\right)\leq\epsilon_{1}$
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\leq 4\alpha^{1+\gamma-\epsilon/2}+\alpha^{1+\gamma/2-\epsilon},$
$-2h(z,v)\leq\eta\leq\frac{1}{2}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+2h(z,v),$
@@ IV-E Achievability
$C_{1}(\alpha)=1+\alpha\log(\alpha)+G_{1}\alpha+\mathcal{O}(\alpha^{3/2-\epsilon}).$ (94)
@@
We follow a similar approach to that of [kanoria2010deletion, Lemma III.1].
Since $\mathbb{X}$ is i.i.d. Bernoulli(1/2) with length $n$, it has a run-length distribution $p_{L}(l)=2^{-l}$. Additionally, the corresponding output $Y(X^{*,n})$ is also i.i.d. Bernoulli with run length $L_{Y}=n+\text{Bin}(n,\alpha)$.
Using the chain rule of entropy, we have
$H(Y(X^{*,n}),L_{Y})$ $=H(L_{Y})+H(Y(X^{*,n})|L_{Y})$ (95)
$=H(L_{Y}|Y(X^{*,n}))+H(Y(X^{*,n})).$ (96)
Since $H(L_{Y}|Y(X^{*,n}))=0$, we have
$H(Y(X^{*,n}))=H(L_{Y})+H(Y(X^{*,n})|L_{Y}).$ (97)
Note that for $H(Y(X^{*,n})\mid L_{Y})$, we have
$H(Y(X^{*,n})|L)$ $=\sum_{l}H(Y(X^{*,n})|L=l)p_{L_{Y}}(l)$
$=\sum_{l}lp_{L_{Y}}(l).$ (98)
Since $L_{Y}=n+\text{Bin}(n,\alpha)$, one can equivalently write
$H(Y(X^{*,n})|L_{Y})=n(1+\alpha).$ (99)
Furthermore, since $L_{Y}$ has shifted binomial distribution, its entropy is $\mathcal{O}(\log(n))$. Hence, one can conclude that
$H(Y(X^{*,n}))$ $=n(1+\alpha)+\mathcal{O}(\log(n)).$ (100)
@@ IV-F Converse
The proof follows a similar structure to that of [kanoria2010deletion, Lemma III.2] with some modifications.
@@
We further define $F=F(\mathbb{X},\mathbb{A})$ as a binary vector with the same length as the channel output. The elements of $F$ are 1 wherever the corresponding bit in $Y_{L^{*}}$ is flipped relative to $Y$, and 0 otherwise. The entropy $H(F)$ can be calculated by:
@@
where $L_{F}$ is the length of $F$. The rest of the proof follows the same steps as [kanoria2010deletion, Lemma III.2]. ∎
@@
For any $\epsilon>0$ there exists there exists $\alpha_{0}=\alpha_{0}(\epsilon)>0$ such that the following happens. For any $L^{*}\in\mathbb{N}$ and any $\mathbb{X}\in\mathcal{S}_{L^{*}}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$. If $\alpha<\alpha_{0}(\epsilon)$, then
$I(\mathbb{X})\leq 1+\alpha\log\alpha+G_{1}\alpha+\alpha^{3/2-\epsilon}.$ (107)
@@
With a similar approach to that of [kanoria2010deletion, Lemma III.3], we can write
@@
since $Y(X^{n})$ contains $n+\text{Binomial}(n,\alpha)$ bits. The proof involves using the lower bound in Corollary 1 along with $h(z,v)$.
Note that $h(z,v)=h(z)+h(v|z)\leq h(z)+h(v)$. Recall the definitions of $\mathbb{Z}$ and $\mathbb{V}$ given in (21) and (22), respectively: $\mathbb{Z}$ is the process used to keep track of reversed insertions, while $\mathbb{V}$ stores the inserted bits in the corresponding positions. Note that, due to the relationship between $\mathbb{Z}$ and $\mathbb{V}$, we have $h(v)\leq h(z)$ which results in $h(z,v)\leq 2h(z)$.
With these definitions, the expected number of reversed insertions in a run of length $l$ can be upper bounded by
$l\alpha-l\alpha(1-\alpha)^{l-1}\leq l^{2}\alpha^{2},$
where the inequality follows from the fact that $(1-\alpha)^{l-1}\geq 1-(l-1)\alpha$.
Since a run can have at least length one, we have $z=P(Z_{i}=1)\leq\alpha^{2}\mathbb{E}[L^{2}]$.
Using [kanoria2013optimal, Lemma V.3] and ${L^{*}}\leq 1/\alpha$ with probability 1, we have
$\mathbb{E}[{L}^{2}]\leq\frac{1}{\alpha}E[L]\leq 20\alpha^{\gamma-1},$ (109)
which implies $z\leq 20\alpha^{\gamma+1}$. It follows that
$h({z})\leq\alpha^{1+\gamma-\epsilon/2},$ (110)
for small enough $\alpha$.
Using $h(z,v)\leq 2h(z)$, we have
$h({z},v)\leq 2\alpha^{1+\gamma-\epsilon/2}.$ (111)
Also noting
@@ V Capacity Approximation for the Gallager Insertion Channel
Let $C_{2}(\alpha)$ be the capacity of the Gallager insertion channel with insertion probability $\alpha$. Then, for small $\alpha$ and any $\epsilon>0$, we have
$C_{2}(\alpha)=1+\alpha\log(\alpha)+G_{2}\alpha+\mathcal{O}(\alpha^{3/2-\epsilon}),$ (113)
@@
$+\frac{1}{4}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}(a+b+2)2^{-a}2^{-b}h\left(\frac{a+1}{a+b+2}\right),$ (114)
@@ V-A Calculation of $H(A^{n},B^{n})$
$P(a,b^{1},b^{2})=\begin{cases}\frac{\alpha}{4},&\text{if }(a,b^{1},b^{2})\in\{(1,1,0),(1,0,1),\\ &\hskip 18.49988pt\hskip 18.49988pt(1,1,1),(1,0,0)\}\\ 1-\alpha,&\text{if }(a,b^{1},b^{2})=(0,0,0)\\ 0,&\text{otherwise}\end{cases}$
@@
As in Section IV-B, we introduce a modified insertion process for the Gallager insertion channel by defining $\hat{A}^{n}$ and $\hat{B}^{n}=(\hat{B}^{1,n},\hat{B}^{2,n})$. The modified process keeps the runs with at most one insertion as they are while reversing more insertions.
@@
$P$ $(z=a,v^{1}=b,v^{2}=c)=\begin{cases}&1-\alpha+\mathbb{E}[\alpha(1-\alpha)^{L_{0}-1}],\\ &\hskip 18.49988pt\text{if }(a,b,c)=(0,0,0)\\ &\frac{1}{4}\left(\alpha-\mathbb{E}[\alpha(1-\alpha)^{L_{0}-1}]\right),\\ &\hskip 18.49988pt\text{if }(a,b,c)\in\mathcal{A}\\ &0,\hskip 18.49988pt\text{otherwise},\end{cases}$ (124)
@@
As in Sec. IV-B, one can use (124) to calculate $H(Z^{n},{V}^{1,n},{V}^{2,n})\leq nh(z,v^{1},v^{2})$ and show that it does not affect lower order terms.
@@
Consider the Gallager insertion channel model with any $\mathbb{X}$ such that $\mathbb{E}[L_{0}\log L_{0}]{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{<\infty}}$. Then, we have
@@
where $2h(z,v^{1},v^{2})\leq\eta\leq\frac{1}{4}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+\frac{1}{4}\alpha^{2}(\mathbb{E}[L_{0}]-1)+2h(z,v^{1},v^{2})$.
@@
where $0\leq\delta\leq\frac{1}{4}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+\frac{1}{4}\alpha^{2}(\mathbb{E}[L_{0}]-1)$.
@@
[…]
@@
[…]
@@
[…]
@@
$P(\hat{A}_{i}=1)\in(\alpha-l_{(i)}\alpha^{2},\alpha).$ (137)
@@
where $0\leq\delta\leq\frac{1}{4}\alpha^{2}\mathbb{E}[L_{0}\log(L_{0})]+\frac{1}{4}\alpha^{2}(\mathbb{E}[L_{0}]-1)$, which completes the proof of Lemma 16. ∎
@@
For the Gallager insertion channel any $\epsilon>0$, there exists $\alpha_{0}\equiv\alpha_{0}(\epsilon)>0$ such that for any $\alpha<\alpha_{0}$ and any stationary and ergodic input process $\mathbb{X}\in\mathcal{S}_{\lfloor 1/\alpha\rfloor}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$, we have:
$\lim_{n\rightarrow\infty}\frac{1}{n}H(K|X^{n},Y)=$ $\frac{\alpha}{4}\sum_{a=1}^{\infty}\sum_{b=1}^{\infty}(a+b+2)2^{-a}2^{-b}h\left(\frac{a+1}{a+b+2}\right)+\epsilon_{1},$ (139)
@@
$\alpha^{1+\gamma/2-\epsilon}-4\alpha^{1+\gamma-\epsilon/2}-\frac{\alpha^{2}}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1})(r_{j}+r_{j+1}+2)h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right)\leq\epsilon_{1}\leq 4\alpha^{1+\gamma-\epsilon/2}+\alpha^{1+\gamma/2-\epsilon},$ (140)
@@
[…]
@@
For the Gallager insertion channel using the perturbed insertion process with any $\epsilon>0$, there exists $\alpha_{0}\equiv\alpha_{0}(\epsilon)>0$ such that for any $\alpha<\alpha_{0}$ and any stationary and ergodic input process $\mathbb{X}\in\mathcal{S}_{\lfloor 1/\alpha\rfloor}$ satisfying $H(\mathbb{X})>1-\alpha^{\gamma}$, where $\gamma>\frac{1}{2}$, we have:
$\lim_{n\rightarrow\infty}H(\check{K}|X^{n},\check{Y})=$ $\frac{\alpha}{4}\sum_{r_{j}=1}^{\infty}\sum_{r_{j+1}=1}^{\infty}(r_{j}+r_{j+1}+2)2^{-r_{j}}2^{-r_{j+1}}h\left(\frac{r_{j}+1}{r_{j}+r_{j+1}+2}\right)-\epsilon_{2}+\epsilon_{3}$ (141)
@@
$\alpha^{1+\gamma/2-\epsilon}\leq\epsilon_{3}\leq\alpha^{1+\gamma/2-\epsilon}.$ (143)